# Rethinking Diffusion for Text-Driven Human Motion Generation

### v1

@@ Abstract
Since 2023, Vector Quantization (VQ)-based discrete generation methods have rapidly dominated human motion generation, primarily surpassing diffusion-based continuous generation methods in standard performance metrics. However, VQ-based methods have inherent limitations. Representing continuous motion data as limited discrete tokens leads to inevitable information loss, reduces the diversity of generated motions, and restricts their ability to function effectively as motion priors or generation guidance. In contrast, the continuous space generation nature of diffusion-based methods makes them well-suited to address these limitations and with even potential for model scalability. In this work, we systematically investigate why current VQ-based methods perform well and explore the limitations of existing diffusion-based methods from the perspective of motion data representation and distribution. Drawing on these insights, we preserve the inherent strengths of a diffusion-based human motion generation model and gradually optimize it with inspiration from VQ-based approaches. Our approach introduces a human motion diffusion model enabled to perform bidirectional masked autoregression, optimized with a reformed data representation and distribution. Additionally, we also propose more robust evaluation methods to fairly assess different-based methods. Extensive experiments on benchmark human motion generation datasets demonstrate that our method excels previous methods and achieves state-of-the-art performances.
@@ 1 Introduction
Figure 1: HumanML3D FID results. Each bubble’s area is proportional to the model size. We achieve superior performance and demonstrate model scalability with a diffusion-based approach.
Despite the improvement in performance, VQ-based methods still exhibit notable limitations. Representing continuous motion data as limited groups of discrete tokens inherently causes a loss of motion information, reduces generation diversity, and limits their ability to serve as motion priors or generation guidance (e.g., in the generation of dual-human motion and human-object interactions). Moreover, unlike language models, these discrete tokens often lack sufficient contextual richness affecting model scalability.
In contrast, the continuous space nature of diffusion-based generation methods can effectively address these limitations and offers potential for model scalability demonstrated by many image diffusion models [28, 22], making it an appealing alternative. In light of these, researchers have started to revisit diffusion-based approaches [47, 45]. However, these attempts struggle to achieve comparable performance to VQ-based methods. More importantly, the reasons for the performance gap between VQ and diffusion-based human motion generation methods remain unclear.
In this work, we first systematically investigate why VQ-based motion generation approaches perform well and explore the limitations of diffusion-based methods from the perspective of motion data representation and distribution. Specifically, we: (1) Examine how VQ-based discrete formulations benefit from training with current motion data representations consisting of redundant dimensions while existing diffusion models are hindered by this data composition and its distribution. (2) Explore how VQ-based methods inherently align with current evaluation metrics, which incorporate the entire data representation including redundant dimensions, whereas diffusion-based methods are often penalized under these evaluation criteria.
Based on our diagnostic findings and inspirations from current VQ-based approaches, we aim to close the performance gap by gradually enhancing a diffusion-based model tailored for human motion generation. We first restructure the motion data representation by excluding redundant information and processing it through a 1D ResNet [15]-based AutoEncoder to achieve a distribution more compatible with diffusion models. This restructuring also improves per-frame smoothness and opens possibilities for diffusion genres beyond DDPM [18]. Additionally, we build a diffusion-based motion generation model with bidirectional masked autoregressive strategies. Finally, we propose more robust evaluators that provide unbiased assessments of the different approaches.
@@
  • We systematically compare and investigate the reasons why VQ-based methods outperform diffusion-based methods from motion data representation and distribution perspective, providing analysis with theoretical and experimental support.
  • Based on our findings and VQ-based model inspirations, we propose a scalable bidirectional masked autoregressive diffusion-based generation framework that leverages reformed data representation and distribution, evaluated with more robust evaluation methods.
  • Our method achieves new state-of-the-art performance on text-to-motion generation task, with significant improvements on both KIT-ML [32] and HumanML3D [11] datasets.
@@ 2 Diagnosis: Motion Data Representation and Distribution
In this section, we systematically analyze how motion representation and distribution impact training, sampling (Sec. 2.2), and evaluation robustness (Sec. 2.3), revealing why these factors favor VQ-based methods while limiting diffusion-based approaches.
@@ 2.1 Prelimenary
VQ-based Human Motion Generation Methods. The VQ-based methods primarily adopt two types of Vector Quantized-Variational AutoEncoders (VQ-VAEs)[41]: a standard VQ-VAE[43] or residual (R) VQ-VAE [13], to transform motion data into discrete tokens.
Given a motion sequence $\mathbf{x}^{1:N}\in\mathbb{R}^{N\times D}$of length $N$, the transformation begins by encoding $\mathbf{x}^{1:N}$ into a latent sequence $\mathbf{h}^{1:n}\in\mathbb{R}^{n\times d}$ with a 1D convolutional encoder $\mathrm{E}$. For vanilla VQ-VAE, each vector is quantized via a base layer VQ codebook to the nearest token. In RVQ-VAE, $V$ additional residual quantization layers are used with each residual layer quantifying the difference between the original latent vector and the quantized representation from the preceding layers. The indices $k$ from the base layer (VQ-VAE) or all layers (RVQ-VAE) form discrete inputs for training generative models with sequence generation.
The generated discrete sequence $\mathbf{g}=\mathbf{g}_{0:n}$ (or $\mathbf{g}_{0:n}^{0:V+1}$ in residual case) is embedded via the codebook, then projected back (or summed then project in residual case) with an upsample convolutional decoder $\mathrm{D}$ to obtain the final motion.
@@
During sampling, starting from random noise $\mathbf{x}_{T}\sim\mathcal{N}(0,\mathbf{I})$, for each $t$, the model with $\mathbf{x}_{t}$ as input, predicts the original motion $\mathbf{x}_{0}$ or the noise $\boldsymbol{\epsilon}$ (then to $\mathbf{x}_{0}$ with Eq. 1), and deduces intermediate samples $\mathbf{x}_{t-1}$ (for $t\in{1\text{ to }T}$) as:
$\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\mathbf{x}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})}{1-\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t}\vspace{-1.25mm}$ (2)
@@
Motion Data Representation. The majority of recent methods utilize the canonical pose representation introduced by [11] on widely-used datasets, including KIT-ML [32] and HumanML3D [11]. This representation at a given time position $i$ is defined as $\mathbf{x}^{i}=[\dot{r}^{a},\dot{r}^{xz},\dot{r}^{h},j^{p},j^{v},j^{r},c^{f}]$, comprising seven feature components: root angular velocity $\dot{r}^{a}$, root linear velocities $\dot{r}^{xz}$ in the XZ-plane, root height $\dot{r}^{h}$, local joint positions $j^{p}\in\mathbb{R}^{3(N_{j}-1)}$, local velocities $j^{v}\in\mathbb{R}^{3(N_{j}-1)}$, joint rotations $j^{r}\in\mathbb{R}^{6(N_{j}-1)}$ in local space, and binary foot-ground contact features $c^{f}\in\mathbb{R}^{4}$, where $N_{j}$ denotes the joint number. However, only the first 4 feature groups from this over-parameterized representation are used to produce final human motion, making the remaining 3 components redundant.
@@ 2.2 Impacts on Training and Sampling
Table 1: Impact of Redundant Features on VQ-based Models. VQ-based methods, T2M-GPT and Momask, trained with redundant features exhibit better reconstruction performance and lead to better generation quality on HumanML3D dataset.
@@
Benefiting VQ-Based Methods. The redundancy in data representation benefits VQ-VAE training which then enhances discrete generation modeling. To validate this viewpoint, we conduct controlled experiments by training VQ-VAE from T2M-GPT [43] and RVQ-VAE from MoMask [13] on HumanML3D with and without redundancy, and then train each generative models following original methods. As shown in Tab. 1, training with redundant dimensions results in significantly better VQ-VAEs which subsequently enhance the performance of discrete generative models. To understand the reason, we further analyze their role in VQ-VAE training.
Let $\mathbf{x}^{\text{GT}}_{\text{r}}$ and $\mathbf{x}^{\text{GT}}_{\text{e}}$ denote the ground truth data with and without redundancy and $\mathbf{x}^{\text{pred}}_{\text{r}}$ and $\mathbf{x}^{\text{pred}}_{\text{e}}$ represent the prediction. The reconstruction loss $\mathcal{L}^{\text{rec}}_{r}$ and $\mathcal{L}^{\text{rec}}_{e}$ with and without redundancy, typically measured with variation of $L_{1}$ or $L_{2}$ loss can be decomposed as:
@@
$=L^{\text{rec}}_{e}+\mathcal{L}(\mathbf{x}^{\text{GT}}_{\text{r-e}}-\mathbf{x}^{\text{pred}}_{\text{r-e}}).\vspace{-1.5mm}$
showing the redundancy acts as a data-level regularization
This built-in regularization helps reduce model variance, producing representations less sensitive to fluctuations in training data, leading to better generalization which ultimately enhances codebook usage and consistency. As shown in Fig. 3 on the HumanML3D test set, both VQ-VAE and RVQ-VAE and models trained with redundant dimensions exhibit more uniform code utilization while models trained without redundancy show distinct spikes in certain codes meaning overfitting, demonstrating the benefit of incorporating redundant dimensions for VQ-based methods.
Table 2: The results of MDM on humanML3D dataset. We report the results of MDM with original $\mathbf{x}_{0}$ prediction vs. with $\boldsymbol{\epsilon}$ prediction. Training to predict $\mathbf{x}_{0}$ leads to significant better results
@@
MDM-50Step [39] $\boldsymbol{\epsilon}$ $31.265$ $0.054$ $0.103$ $0.147$
@@
Most diffusion-based methods follow the DDPM [18] but predict original motion $\mathbf{x}_{0}$. Attempts to train these methods to predict noise often fail, e.g. MDM [39] in Tab. 2. This deterministic $\mathbf{x}_{0}$-only prediction in each timestep also limits training simplicity and sampling diversity.
Below, we outline how current motion data constrains diffusions to learn and sample only $\mathbf{x}_{0}$: dimensional distribution mismatch and error amplification with $\boldsymbol{\epsilon}$ prediction.
@@
First, the current motion data may not follow a standard normal distribution with standard z-normalization due to its mixed structure consisting of features from 3D continuous, 6D rotatory (e.g. joint rotation), and categorical (e.g. foot contacts) distribution. This leads to a dimensional distribution mismatch and challenges the interpolation function (Eq. 1 for DDPM). In the forward diffusion process, due to the differing initial distributions of $\mathbf{x}_{0}$’s different feature groups, feature groups of $\mathbf{x}_{T}$ may converge to their own distinct distribution by time step $T$, rather than all converging to a standard normal distribution. Consequently, in the reverse diffusion process, starting from a standard normal distribution, which the model is not trained to predict from, leads to errors in motion generation.
@@
Second, consider two main data processing methods: (1) standard z-normalization, followed by averaging the standard deviation (SD) $\sigma^{\mathbf{x}}$ for each 7 feature groups, (2) additionally, scaling the SD by a feature bias term. In the first method, the averaged SD $\sigma^{{}^{\prime}\mathbf{x}}$ for a group with $D$ dimensions is: $\sigma^{{}^{\prime}\mathbf{x}}=\frac{\sum_{i=0}^{D-1}\sigma^{\mathbf{x}}_{i}}{D}$ with the ratio $\phi^{\prime\mathbf{x}}$ of original (which is also $\boldsymbol{\epsilon}$’s SD = $\mathbf{I}$) to adjusted SD as $\phi^{\prime\mathbf{x}}=\frac{\sigma^{\mathbf{x}\prime}}{\sigma^{\mathbf{x}}}$. In the second method, this ratio is further adjusted by a feature bias term $\gamma$: $\phi^{\prime\mathbf{x}}=\gamma\times\frac{\sigma^{\mathbf{x}\prime}}{\sigma^{\mathbf{x}}}$.
Unlike predicting $\mathbf{x}_{0}$, where the error is standardized to predicted $\mathbf{x}_{0}$ and its ground truth in each timestep, this SD ratio will cause error amplification when predicting noise.
@@
Suppose $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ denote the predicted noise $\boldsymbol{\epsilon}$ at time $t$ and define $\delta_{\mathbf{x}_{0}}$ as the squared error between ground truth and predicted $\mathbf{x}_{0}$, and $\delta_{\boldsymbol{\epsilon}}$ as the error for noise prediction. Then we have $\delta_{\boldsymbol{\epsilon}}=\|\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\boldsymbol{\epsilon}\|_{2}^{2}$ and $\delta_{\mathbf{x}_{0}}=\left\|\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)-\mathbf{x}_{0}\right\|_{2}^{2}$. Substitute $\mathbf{x}_{0}$ from Eq. 1 (detailed deduction in App. A.1), we get:
@@
$\delta_{\mathbf{x}_{0}}=\left\|\frac{1-\bar{\alpha}_{t}}{\bar{\alpha}_{t}}\right\|^{2}_{2}\delta_{\boldsymbol{\epsilon}},\vspace{-1.5mm}$ (4)
@@
an standard error relation $\delta_{\boldsymbol{\epsilon}}\to\delta_{\mathbf{x}_{0}}$ if $\mathbf{x}_{0}$ is processed correctly which only responds to time coefficient $\bar{\alpha}$. Since the SD ratio $\phi^{\prime\mathbf{x}}$ applies to both deduced and ground truth $\mathbf{x}_{0}$, then Eq. 4 updates to:
@@
$\delta_{\mathbf{x}_{0}}\times\phi^{\prime}_{i}$ $=\left\|\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}\right\|_{2}^{2}\delta_{\boldsymbol{\epsilon}}.\vspace{-2.25mm}$ (5)

### v2

@@ Abstract
Since 2023, Vector Quantization (VQ)-based discrete generation methods have rapidly dominated human motion generation, primarily surpassing diffusion-based continuous generation methods in standard performance metrics. However, VQ-based methods have inherent limitations. Representing continuous motion data as limited discrete tokens leads to inevitable information loss, reduces the diversity of generated motions, and restricts their ability to function effectively as motion priors or generation guidance. In contrast, the continuous space generation nature of diffusion-based methods makes them well-suited to address these limitations and with even potential for model scalability. In this work, we systematically investigate why current VQ-based methods perform well and explore the limitations of existing diffusion-based methods from the perspective of motion data representation and distribution. Drawing on these insights, we preserve the inherent strengths of a diffusion-based human motion generation model and gradually optimize it with inspiration from VQ-based approaches. Our approach introduces a human motion diffusion model enabled to perform masked autoregression, optimized with a reformed data representation and distribution. Additionally, we propose a more robust evaluation method to assess different approaches. Extensive experiments on various datasets demonstrate our method outperforms previous methods and achieves state-of-the-art performances.
@@ 1 Introduction
Figure 1: The FID results on HumanML3D dataset. The bubble size is proportional to the model size. We achieve superior performance and demonstrate model scalability.
Despite the performance improvement, VQ-based methods still exhibit notable limitations. Representing continuous motion data as limited groups of discrete tokens inherently causes a loss of motion information, reduces generation diversity, and limits their ability to serve as motion priors or generation guidance (e.g., in the generation of dual-human motion and human-object interactions). Moreover, unlike language models, these discrete tokens often lack contextual richness, which can hinder model scalability.
In contrast, the continuous space nature of diffusion-based generation methods effectively overcomes these limitations and offers potential for model scalability, as evidenced by numerous image diffusion models [66, 51]. In light of these, researchers have started to revisit diffusion-based approaches [117, 114]. However, these attempts struggle to achieve performance comparable to that of VQ-based methods. More importantly, the reasons for the performance gap between VQ and diffusion-based human motion generation methods remain unclear.
In this work, we first systematically investigate why VQ-based motion generation approaches perform well and explore the limitations of diffusion-based methods from the perspective of motion data representation and distribution. Specifically, (1) We examine how VQ-based discrete formulations benefit from training with current motion data representation, which consists of redundant dimensions, while this data composition and its distribution hinder existing diffusion models. (2) We explore how VQ-based methods inherently align with current evaluation metrics, which incorporate the entire data representation including redundant dimensions, whereas diffusion-based methods are often penalized under these evaluation criteria.
Based on our diagnostic findings and inspirations from current VQ-based approaches, we aim to close the performance gap by gradually enhancing a diffusion-based model tailored for human motion generation. We first restructure the motion data representation by removing redundant information, resulting in a distribution better suited for diffusion models. We then process them with a 1D ResNet [29]-based AutoEncoder for better per-frame smoothness. Additionally, we build a diffusion-based motion generation model with masked autoregressive strategies to simplify training objective. Finally, we propose more robust evaluators for unbiased assessments of different approaches.
@@
  • We systematically investigate the reasons why VQ-based methods outperform diffusion-based methods from motion data representation and distribution perspective, providing analysis with theoretical and experimental support.
  • Inspired by our diagnostic findings, we propose a scalable masked autoregressive diffusion-based generation framework and a more robust evaluation method.
  • Our method achieves state-of-the-art performance on text-to-motion generation task with significant improvements on KIT-ML [77] and HumanML3D [24] datasets.
@@ 2 Diagnosis: Motion Data Representation and Distribution
In this section, we study how motion representation and distribution impact training, sampling (Sec. 2.2), and evaluation robustness (Sec. 2.3), revealing how these factors favor VQ-based methods but limit diffusion-based approaches.
@@ 2.1 Prelimenary
VQ-based Human Motion Generation Methods primarily adopt either a standard Vector Quantized Variational AutoEncoders (VQ-VAEs) [92, 111] or a residual (R) VQ-VAE [26] to map motion data into discrete tokens.
Given a motion sequence $\mathbf{x}^{1:N}\in\mathbb{R}^{N\times D}$ of length $N$ and human pose dimension $D$, the transformation begins by encoding $\mathbf{x}^{1:N}$ into a latent sequence $\mathbf{h}^{1:n}\in\mathbb{R}^{n\times d}$ with a 1D convolutional encoder $\boldsymbol{\mathrm{E}}$. For vanilla VQ-VAE, each vector is quantized via a base VQ codebook to the nearest token. In RVQ-VAE, additional residual quantization layers are used to quantify the difference between the original latent vector and the quantized representation from the preceding layers. The indices $\mathbf{k}$ from the base layer (VQ-VAE) or all layers (RVQ-VAE) form discrete inputs for training generative models with sequence generation.
The generated discrete sequence $\mathbf{g}=\mathbf{g}_{0:n}$ (or $\mathbf{g}_{0:n}^{0:V+1}$ in residual case) is embedded via the codebook, then projected back (or summed then project in residual case) with an upsample convolutional decoder $\boldsymbol{\mathrm{D}}$ to obtain the final motion.
@@
During sampling, starting from random noise $\mathbf{x}_{T}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, for each $t$, the model with $\mathbf{x}_{t}$ as input, predicts the original motion $\mathbf{x}_{0}$ or the noise $\boldsymbol{\epsilon}$ (then to $\mathbf{x}_{0}$ with Eq. 1), and deduces intermediate samples $\mathbf{x}_{t-1}$ (for $t\in{1\text{ to }T}$):
$\frac{\sqrt{\alpha_{t}}(1-\bar{\alpha}_{t-1})}{1-\bar{\alpha}_{t}}\mathbf{x}_{t}+\frac{\sqrt{\bar{\alpha}_{t-1}}(1-\alpha_{t})}{1-\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t}\vspace{-2mm}$ (2)
@@
Motion Data Representation. The majority of recent methods utilize the canonical pose representation introduced by [24] on widely-used datasets, including KIT-ML [77] and HumanML3D [24]. This representation at a given time step $i$ is defined as $\mathbf{x}^{i}=[\dot{r}^{a},\dot{r}^{xz},\dot{r}^{h},j^{p},j^{v},j^{r},c^{f}]$, comprising seven feature components: root angular velocity $\dot{r}^{a}$, root linear velocities $\dot{r}^{xz}$ in the XZ-plane, root height $\dot{r}^{h}$, local joint positions $j^{p}\in\mathbb{R}^{3(N_{j}-1)}$, local velocities $j^{v}\in\mathbb{R}^{3(N_{j}-1)}$, joint rotations $j^{r}\in\mathbb{R}^{6(N_{j}-1)}$ in local space, and binary foot-ground contact features $c^{f}\in\mathbb{R}^{4}$, where $N_{j}$ denotes the joint number. However, only the first 4 feature groups from this over-parameterized representation are used to produce final human motion, making the remaining 3 components redundant. We categorize the first 4 feature groups as essential features while classifying the remaining as redundant.
@@ 2.2 Impacts on Training and Sampling
Table 1: Impact of redundant features on VQ-based models. VQ-based methods, T2M-GPT and MoMask, trained with redundant features exhibit better reconstruction performance and lead to better generation quality on the HumanML3D dataset.
@@
Benefiting VQ-Based Methods. The redundancy in data representation benefits VQ-VAE training which then enhances discrete generative modeling. To validate this viewpoint, we conduct ablative experiments by training VQ-VAE from T2M-GPT [111] and RVQ-VAE from MoMask [26] on HumanML3D, comparing their reconstruction performance with and without redundant data representations. Then we train T2M-GPT and MoMask following their original methods, both with and without redundant data representations, and evaluate their performance using human motion generation metrics. As shown in Tab. 1, training with redundant dimensions results in VQ-VAEs with significantly better reconstruction performance, which subsequently enhances the performance of discrete generative models. To understand the reason, we further analyze the role of data representation in VQ-VAE training.
Let $\mathbf{x}^{\text{GT}}_{\text{r}}$ and $\mathbf{x}^{\text{GT}}_{\text{e}}$ denote the ground truth data with and without redundancy and $\mathbf{x}^{\text{pred}}_{\text{r}}$ and $\mathbf{x}^{\text{pred}}_{\text{e}}$ represent the prediction. The reconstruction loss $\mathcal{L}^{\text{rec}}_{r}$ with redundancy can be decomposed as:
@@
$=L^{\text{rec}}_{e}+\mathcal{L}(\mathbf{x}^{\text{GT}}_{\text{r-e}}-\mathbf{x}^{\text{pred}}_{\text{r-e}}).\vspace{-2mm}$
where $\mathcal{L}^{\text{rec}}_{e}$ is the reconstruction loss without redundancy representation, and $\mathcal{L}$ is typically measured with variation of $L1$ or $L2$ loss. This decomposition shows the redundancy acts as a data-level regularization
This built-in regularization reduces model variance, making representations less sensitive to training data fluctuations. This improvement enhances generalization, ultimately leading to more effective and consistent codebook usage, as shown in Fig. 3. Specifically, on the HumanML3D test set, both VQ-VAE and RVQ-VAE, which were trained on the data with redundant dimensions, exhibit more uniform code utilization. In contrast, models trained on the data without redundancy show distinct spikes in certain codes. Numerically, for VQ-VAE, the KL divergence against uniform distribution is 0.205 vs. 0.337, and for RVQ-VAE, it is 0.243 vs. 0.351, using original vs. essential dimensions. These results demonstrate the benefit of incorporating redundancy for VQ-based methods.
Table 2: The results of MDM on humanML3D dataset. We report the results of MDM with original $\mathbf{x}_{0}$ prediction vs. with $\boldsymbol{\epsilon}$ prediction. Training to predict $\mathbf{x}_{0}$ leads to better results.
@@
MDM-50Step [88]-Cosine $\boldsymbol{\epsilon}$ $31.265$ $0.054$ $0.103$ $0.147$
MDM-50Step [88]-Linear $\boldsymbol{\epsilon}$ $1.574$ $0.279$ $0.336$ $0.415$
@@
Most diffusion-based methods follow the DDPM [32] but predict original motion $\mathbf{x}_{0}$. However, this deterministic $\mathbf{x}_{0}$-only prediction in each timestep can limit training simplicity and sampling diversity. Attempts to train these methods to predict noise often result in inaccurate motion, human shape e.g. MDM [88] with cosine beta schedule, or persistent shaking e.g. MDM with linear beta schedule in Tab. 2. Below, we explain the two major factors in detail: dimensional distribution mismatch and error amplification.
@@
First, the current motion data do not follow a standard normal distribution with standard z-normalization due to its mixed structure consisting of features from 3D continuous (e.g. joint position), 6D rotatory (e.g. joint rotation), and categorical (e.g. foot contacts) distribution. This leads to a dimensional distribution mismatch and challenges the interpolation function (Eq. 1 for DDPM). In the forward diffusion process, due to the varying initial distributions across different feature groups of $\mathbf{x}_{0}$, feature groups of $\mathbf{x}_{T}$ may converge to their own distinct distribution by time step $T$, rather than all converging to a same standard distribution. Consequently, in reverse diffusion process, starting from a standard distribution leads to errors in motion generation.
@@
Second, when normalizing the human motion data, the standard deviation (SD) is averaged in each of 7 feature groups as $\sigma^{{}^{\prime}\mathbf{x}}=\frac{\sum_{i=0}^{D-1}\sigma^{\mathbf{x}}_{i}}{D}$, where D is the number of dimensions in each group. We define the ratio $\phi^{\prime\mathbf{x}}=\frac{\sigma^{\mathbf{x}\prime}}{\sigma^{\mathbf{x}}}$, where $\sigma^{\mathbf{x}}$ is the original SD. Then this ratio is further adjusted by a feature bias term $\gamma$: $\phi^{\prime\mathbf{x}}=\gamma\times\frac{\sigma^{\mathbf{x}\prime}}{\sigma^{\mathbf{x}}}$ when the data is fed in the network. This SD ratio will cause error amplification when predicting noise.
@@
Suppose $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ denote the predicted noise $\boldsymbol{\epsilon}$ at time $t$ and define the squared error $\delta_{\mathbf{x}_{0}}$ between ground truth and predicted $\mathbf{x}_{0}$, and error $\delta_{\boldsymbol{\epsilon}}$ for noise prediction. Then we have $\delta_{\boldsymbol{\epsilon}}=\|\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)-\boldsymbol{\epsilon}\|_{2}^{2}$ and $\delta_{\mathbf{x}_{0}}=\left\|\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)-\mathbf{x}_{0}\right\|_{2}^{2}$. Substitute $\mathbf{x}_{0}$ from Eq. 1 (detailed deduction in App. A.1), we get:
@@
$\delta_{\mathbf{x}_{0}}=\left\|\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}\right\|^{2}_{2}\delta_{\boldsymbol{\epsilon}},\vspace{-1.5mm}$ (4)
@@
an standard error relation $\delta_{\boldsymbol{\epsilon}}\to\delta_{\mathbf{x}_{0}}$. If $\mathbf{x}_{0}$ is processed correctly, the relation between $\delta_{\mathbf{x}_{0}}$ and $\delta_{\boldsymbol{\epsilon}}$ only responds to time coefficient $\bar{\alpha}$. Since the SD ratio $\phi^{\prime\mathbf{x}}$ applies to both predicted and ground truth $\mathbf{x}_{0}$, then Eq. 4 updates to:
@@
$\delta_{\mathbf{x}_{0}}\times\phi^{\prime}_{i}$ $=\left\|\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}\right\|_{2}^{2}\delta_{\boldsymbol{\epsilon}},\vspace{-2.25mm}$ (5)